when business firms are able to control their basic costs, I'd success assured
Answer #1

Basic costs are direct material cost, direct labor cost, direct expenses, and variable overhead. These costs are crucial for a business to run and the aggregate of these costs constitutes the majority of the total cost. Therefore, a little control over this cost can make a business highly successful, since its profit margin increases.

Example of such cost is “cost of goods sold”, suppose it constitutes 65% of the total cost of $130,000. Therefore, cost of goods sold is [130,000 × 65% =] $84,500. Now COGS decreases to 62%; therefore, COGS = [130,000 × 62% =] $80,600. This makes a saving of cost of [84,500 – 80,600 =] $3,900.